If u = (-1/r)r, find the value of grad (div u) .
1
Expert's answer
2021-08-31T17:09:07-0400

u=(-1/r)r

div u=div[(-1/r)r]=-2/r

grad (div u)=[grad (-2/r)= (2/r3) grad r]

grad f(r)=[f'(r) grad r]

[f'(r) grad r]= 2/r3(1/r)r= (2/r3)r